North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog Saturday 10th March 2018

Saturday 10th March 2018

Ian (who was present following a cancelled diesel course), reports a murky, wet, morning at Grosmont with Andy and Joan Lowes, Peter Ellis, Trevor Wilford, and Bryan Orange and the JV's. No mice to report again but he thinks new traps are required as the chocolate has been consumed from our trap - can Bill bring some new traps please! There was fantastic support from some of the JV's parents, with Ethan Humble's Mum and Grandmother manning a sales stall, with cakes, tea and coffee. Another JV's Dad, Andy Pritchard, set up a small model railway layout alongside the stalls. These all attracted lots of visitors into Deviation Shed and also watching Tornado passing by on its way to Pickering, the B1 shunting around the yard and steam heating stock at the Station, and Standard Tank 80136 on passenger work. About £110 was made at the stalls - a tremendous effort and much appreciated by the Group.

Some JVs were working on the J27, some with Piglet working on Lucie, and some working on the Q6. Both valve bores were cleaned with fine emery paper by one of our JVs, with Joan Lowes supervising other JVs on other work, including full greasing of the locomotive and tender. One new JV was working with Bryan, refitting plugs back on to an extension cable and light fitting which got damaged last week. Peter spent the day making new cotters for the eccentrics, and has made and fitted cotters to the left hand side eccentrics. Some of these will need the locking nut reducing slightly to make a better fit, and hopefully Ian McCall will be able to operate the lathe in our workshop to do this on Wednesday. Trevor removed an old stud which was jammed in the regulator valve and fitted two new studs. This valve has now been put back upstairs on the shelf. Andy, with one of the JV's in attendance, worked on each valve in the workshop, completing the assembly with cotter pins ready for fitting. Ian has emphasisied that if the valves have to be dismantled for any reason, the new keys which Andy fitted inside at the rear end of the valve have a small hole tapped into them. An M4 bolt can be screwed into the key for easy removal. Ian himself was working fitting taper pins into the front end of the eccentric rods: a long slow process getting pins from the shed, cutting them to size and filling and splitting. The top two are still to do on Wednesday.

Lunch was had down at the MPD mess room. Afterwards, Andy, Trevor and JV Henry Pritchard fitted the valves. These are still to connect to the valve rods, but the cotters seem to be too loose when fitted and it looks as if new ones will have to be made: the old ones are on the pit wall below each valve. Two strops were found by Trevor on the engine which seem to have missed the inspection a couple of weeks ago. These are 2metre I ton strops Nos 856525 and 856527 and they have been put upstairs in the metal filling cabinet, 3rd drawer down . If so they will have to be inspected next time there is an inspection about 6 months time but in the meantime do not use these strops until their status has been clarified.

Trevor will not be available for the next three Saturdays, and next Wednesday is the last that Ian McCall will be available for six weeks.

The next working Party will be on Wednesday (14th March), by when, hopefully the last of the flue tube holes will have been chamfered so work on the boiler can be progressed.